Grade 10General Physics

What is the total positive or negative charge on a 3.11g of copper penny. Avogadro no. to be asumed as 6.02 * 10^23.

Jatin chauhan
11 Years agoGrade 10

2 Answers

Yash Jain
11 Years ago
See, the total charge over the neutral Copper penny would be zero. that is,
Qnet = Qe + Qp  {where e and p are electrons and protons resp. ; Qep>0}
Individually, Qand Qp can be calculated as follows:
The total number of electrons in an atom of Cu is 29
Given mass: 3.11 ; Molar mass: 63.5
Therefore, no. of moles of Cu = 3.11/63.5 = 0.049
So, total numbers of electron would be = 0.049*6.023*1023*29 = N(say)
Hence, the toal charge due to N electrons is
Qe = Ne {where e is the fundamental charge 1.6*10-19}
     = 0.049*6.023*1023*29*1.6*10-19
     = 1.37*10C
 
ankit singh
5 Years ago
See, the total charge over the neutral Copper penny would be zero. that is, Q net = Q e + Q p where e and p are electrons and protons resp. ; Q e p >0 Individually, Q e and Q p can be calculated as follows: The total number of electrons in an atom of Cu is 29 Given mass: 3.11 ; Molar mass: 63.5 Therefore, no. of moles of Cu = 3.11/63.5 =...